OpenStudy (anonymous):

Does this serie converges or diverges? S=((-1)^n)(n^(1/n))sin(1/n) n from 1 to infinity

OpenStudy (anonymous):

\[\sum_{1}^{\infty} (-1)^n \sqrt[n]{n} \sin(1/n)\]

OpenStudy (anonymous):

Who's there c:

OpenStudy (anonymous):

??

OpenStudy (experimentx):

i think that will not converge. for alternating series lim n->inf |f(n)| = 0

OpenStudy (experimentx):

for this particular case, lin n->inf = indeterminate

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \sqrt[n]{n} \sin(1/n)=0?? \]

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

So , how can I prove if the limit is indeterminate?

OpenStudy (experimentx):

No ... sin(1/n) is indeterminate ... and http://www.wolframalpha.com/input/?i=lim+n-%3Einfinity+n^%281%2Fn%29

OpenStudy (experimentx):

oh ... sorry. my mistake.

OpenStudy (experimentx):

i thought n->0 inside sin it's zero.

OpenStudy (anonymous):

ok, so limit is 0 right?

OpenStudy (experimentx):

well ... one condition is satisfied. the other condition is |a_n| > |a_n+1|

OpenStudy (anonymous):

Yep

OpenStudy (experimentx):

i guess this means it does not converge. http://www.wolframalpha.com/input/?i=Sum [%28-1%29^n+n^%281%2Fn%29+*+sin%281%2Fn%29%2C+1%2C+infinity] n^n sin(1/n) > (n+1)^(n+1) sin(1/n+1)

OpenStudy (anonymous):

can you copy paste the link again?

OpenStudy (zarkon):

it converges

OpenStudy (anonymous):

Could you explain it?

OpenStudy (zarkon):

I could

OpenStudy (experimentx):

http://tinyurl.com/6psphl4

OpenStudy (zarkon):

show that \(\sin(1/n)\) decreases for all \(n\) and show that \(n^{1/n}\) decreases for all \(n\ge 3\)

OpenStudy (experimentx):

whats up with this imaginary part http://www.wolframalpha.com/input/?i=plot+%28-1%29%5En+n%5E%281%2Fn%29+*+sin%281%2Fn%29+from+1+to+infintiy

OpenStudy (zarkon):

\[n^{1/n}\] is not real for most negative values of n\]

OpenStudy (anonymous):

Ok so as limit -->0 and it's a decreasing sequence then the series converges by Leibniz test right?

OpenStudy (zarkon):

yes...the alternating series test

OpenStudy (anonymous):

Ok, thank you guys!

OpenStudy (experimentx):

Oh i get it ... i had been doing the opposite and getting infinity instead http://www.wolframalpha.com/input/?i=lim+n-%3Einfinity+%28n%2B1%29%5E%281%2F%28n%2B1%29%29%2Fn%5E%281%2Fn%29